OpenStudy (anonymous):

Letters a-j represent consecutive integers on a number line. If 2b+f=13, what is the value of b?

OpenStudy (anonymous):

f=b+1+1+1+1=b+4 2b+b+4=13 3b=13-4=9 b=3